# Characterisations of Sobolev spaces and constant functions over metric spaces

### v2

@@ 8. The upper bound for the Sobolev norm
We finally return to the other direction of the proof of Theorem 1.6. With the results of Sections 6 and 7 available as substitutes of some Euclidean tools, the proof can now be completed by an adaptation of the Euclidean argument of [19]. The following lemma borrows some ideas from the proof of [19, Lemma 3.1], but is actually simpler. This may be partly attributed to the fact that we only prove equivalences of norms, in contrast to the identities achieved in the presence of Euclidean symmetries in [19].
@@ 9. The limit relation
We will need the following density result, which is most likely known. Since most sources deal with the non-homogeneous space $M^{1,p}$ instead of $\dot{M}^{1,p}$ (see e.g. [27, Lemma 10.2.7]), we provide the short proof for completeness.
@@
so this restriction is $2L$-Lipschitz. Hence it can be extended to a $cL$-Lipschitz function $\tilde{f}$ on all of $X$. (For real-valued functions, the classical McShane extension [27, page 99] preserves the Lipschitz constant, so we may take $c=2$; for $\mathbb{C}\simeq\mathbb{R}^{2}$ -valued functions, $c=2\sqrt{2}$ will fo [27, Corollary 4.1.6]. In fact, this extension is possible even for Banach space $V$-valued functions, but then the constant $c$ may also depend on the parameters of the domain space $X$ [27, Theorem 4.1.12].)
@@ 10. Further remarks on the Poincaré inequality
Proposition 10.2 is somewhat in the spirit of [17, Theorem 4.1], which shows the equivalence of
@@
    a variant, where the right-hand side is an abstract functional $a_{f}(B)$ satisfying a list of postulates given in [17, Definition 3.2].
@@
satisfies the abstract postulates, so their “new” Poincaré inequality is just a version of the usual one (essentially, the one of Lemma 2.4(4)). In the other direction, where the usual Poincaré inequality is deduced from an abstract one, a key property of the abstract inequality is [17, Definition 3.2($V_{p}$)], which looks a bit like our macroscopic Poincaré inequality of Theorem 3.4, but involves a family of balls of arbitrary radius on the right-hand side, instead of the balls $B(x,r)$ of fixed radius $r$ as in Theorem 3.4. Hence, in both directions, [17, Theorem 4.1] and Proposition 10.2 are essentially different.
As another detail, we also note that [17, Theorem 4.1] only deals with complete spaces, while the main part of Proposition 10.2 is independent of this assumption.

### v3

@@ 8. The upper bound for the Sobolev norm
We finally return to the other direction of the proof of Theorem 1.1. With the results of Sections 6 and 7 available as substitutes of some Euclidean tools, the proof can now be completed by an adaptation of the Euclidean argument of [23]. The following lemma borrows some ideas from the proof of [23, Lemma 3.1], but is actually simpler. This may be partly attributed to the fact that we only prove equivalences of norms, in contrast to the identities achieved in the presence of Euclidean symmetries in [23].
@@ 9. The limit relation
We will need the following density result, which is most likely known. Since most sources deal with the non-homogeneous space $M^{1,p}$ instead of $\dot{M}^{1,p}$ (see e.g. [32, Lemma 10.2.7]), we provide the short proof for completeness.
@@
so this restriction is $2L$-Lipschitz. Hence it can be extended to a $cL$-Lipschitz function $\tilde{f}$ on all of $X$. (For real-valued functions, the classical McShane extension [32, page 99] preserves the Lipschitz constant, so we may take $c=2$; for $\mathbb{C}\simeq\mathbb{R}^{2}$ -valued functions, $c=2\sqrt{2}$ will fo [32, Corollary 4.1.6]. In fact, this extension is possible even for Banach space $V$-valued functions, but then the constant $c$ may also depend on the parameters of the domain space $X$ [32, Theorem 4.1.12].)
@@ 10. Further remarks on the Poincaré inequality
Proposition 10.2 is somewhat in the spirit of [18, Theorem 4.1], which shows the equivalence of
@@
    a variant, where the right-hand side is an abstract functional $a_{f}(B)$ satisfying a list of postulates given in [18, Definition 3.2].
@@
satisfies the abstract postulates, so their “new” Poincaré inequality is just a version of the usual one (essentially, the one of Lemma 2.6(4)). In the other direction, where the usual Poincaré inequality is deduced from an abstract one, a key property of the abstract inequality is [18, Definition 3.2($V_{p}$)], which looks a bit like our macroscopic Poincaré inequality of Theorem 3.4, but involves a family of balls of arbitrary radius on the right-hand side, instead of the balls $B(x,r)$ of fixed radius $r$ as in Theorem 3.4. Hence, in both directions, [18, Theorem 4.1] and Proposition 10.2 are essentially different.
As another detail, we also note that [18, Theorem 4.1] only deals with complete spaces, while the main part of Proposition 10.2 is independent of this assumption.